OpenStudy (bohotness):

Which three ordered pairs complete the table to give solutions of the equation y = x – 4? A. (–3, –7), (0, –4), and (3, –1) B. (–3, –7), (0, –4), and (3 1) C. (–3, –1), (0, –4), and (3, 1) D. (–3, –1), (0, –4), and (3, –1)

OpenStudy (bohotness):

http://prntscr.com/61u1bk

OpenStudy (anonymous):

This is quite easy to solve, because it is only y = x - 4, we can just plug in y and x. For example, on A: \[-7 = -3 - 4\] This is correct, because -7 = -7 Do you see how I got this?

OpenStudy (anonymous):

For the second part on A: \[-4 = 0 - 4\] This is correct as well, because -4 = -4

OpenStudy (bohotness):

yes

OpenStudy (anonymous):

Awesome! If you need anymore help, ask more questions. Make sure to close this question, and give a medal if you feel I deserve it.

OpenStudy (bohotness):

You do AND THANK YOU

OpenStudy (bohotness):

a?

OpenStudy (bohotness):

is right? am i correct??

OpenStudy (anonymous):

It would appear so, yes. You can check with the others if you want to be sure.

OpenStudy (bohotness):

k